# LowDiff: Efficient Frequent Checkpointing via Low-Cost Differential for High-Performance Distributed Training Systems

### v1

@@ Abstract
This paper proposes LowDiff, an efficient frequent checkpointing framework that reuses compressed gradients, serving as differential checkpoints to reduce cost. Furthermore, LowDiff incorporates a batched gradient write optimization to persist these differentials to storage efficiently. It also dynamically tunes both the checkpoint frequency and the batching size to maximize performance. We further enhance LowDiff with a layer-wise gradient reusing and snapshotting approach and a CPU-based asynchronous persistence strategy, enabling frequent checkpointing without gradient compression. Experiments on various workloads show that LowDiff can achieve checkpointing frequency up to per iteration with less than 3.1% runtime overhead.
@@ I Introduction
Deep learning has demonstrated remarkable capabilities in the fields of computer vision [49, 20], natural language processing [14, 6] and so on. Recently, the advent of large models like ChatGPT [37] and GPT-4 [38] has captivated both academic and industrial attention. To efficiently train the large models, distributed training systems [27, 26, 46, 44] are widely deployed across workers (or nodes) to accelerate the training process, which iterates over a dataset many times (i.e., epochs), each containing multiple iterations. Nevertheless, distributed training often requires many nodes and a long training time, leading to frequent failures [55]. For instance, a Microsoft study reveals that the mean time between failures (MTBF) in multi-tenant GPU clusters can range from minutes to days [23]; similarly, OPT-175B, equipped with 992 NVIDIA A100 GPUs, experienced 112 failures during its 90-day training period—averaging nearly two failures per day [34].
@@
To handle failures, checkpointing, which regularly writes the model state to storage (e.g., disks) for failure recovery, can confine the potential loss of the training progress to the interval between consecutive checkpoints, only requiring the recovery of the remaining progress. Traditional checkpointing saves the model states (including the model parameters and optimizer parameters) at epoch boundaries, but an epoch often runs for hours, so the epoch-level checkpointing often results in hours of GPU computation wasted on average for every failure recovery [36].
@@
Consequently, state-of-the-art checkpointing studies on fast recovery (e.g., Checkfreq [36] and Gemini [54]) aim to increase the frequency of checkpointing from epochs to iterations (we call frequent checkpointing). Nevertheless, the frequent checkpointing generates a considerable number of checkpoints, whose creation and transmission will incur substantial computational and communication costs, and thus hinder the training process significantly. Recent studies on reducing such costs (e.g., Check-N-Run [15]) introduce differential checkpointing (DC for short), which only tracks and checkpoints the modified part of the model to reduce the checkpoint size, into recommendation systems by leveraging its unique sparse features.
@@
However, when deploying DC with high frequency in general distributed training systems, we observe two challenges arising: 1) without the help of sparse features, frequent DC incurs significant computation cost required for compressing the original differentials into smaller ones (i.e., compressed differentials), thus hindering the training process, and 2) frequent DC causes significant training stalls, as frequent checkpoint writes incur substantial transmissions which can not be overlapped with training, thus degrading the training performance (see §III-A for details).
@@
Fortunately, we also have two findings from the gradient compression techniques [7, 18, 35, 9, 50] widely used in distributed training systems for communication efficiency: 1) the compressed gradients can be reused to construct the compressed differentials for general deep-learning models. 2) Using compressed gradients eliminates the data dependency between training and checkpointing, enabling them to run in parallel. Moreover, the size of compressed gradients is significantly smaller than that of compressed differentials (see §III-B for details).
Therefore, to address the above challenges, the corresponding findings inspire us to leverage the compressed gradients in distributed training systems for low-cost differential checkpointing, so as to improve the overall training performance. Our motivation is that the compressed gradients can act as the compressed differentials with no DC computation cost and with a lower DC transmission cost.
@@
In this paper, we propose LowDiff, an efficient frequent checkpointing framework, whose main idea is to reuse the compressed gradients to serve as differential checkpoints, which not only eliminates the computation cost for differential compression (via reusing) but also reduces the transmission cost for DC writes (due to smaller-size gradients than differentials). Further, to enhance LowDiff, we accelerate the checkpoint writes via gradient batching and optimize the checkpointing performance by tuning the checkpoint frequency and batching size.
@@
  • We design LowDiff, which realizes the reuse of compressed gradients via a queue structure and facilitates zero-copy transmission between the training and checkpointing processes. Atop LowDiff, we propose a batched gradient writing optimization by merging compressed gradients in CPU memory before persisting to storage. We also design a checkpointing configuration optimization to minimize the wasted GPU time (§IV).
  • We further propose LowDiff+ for efficient gradient reuse without compression. By exploiting the layer-by-layer generation of gradients during training, we design a layer-wise gradient reusing and snapshotting approach. We also reconstruct a model state in CPU memory, performing high-frequency in-memory checkpointing and asynchronous persistence (§V).
  • We implement LowDiff atop DeepSpeed, open-sourced at https://github.com/YuchongHu/LowDiff. Experiments on various models show that LowDiff can achieve checkpointing as frequently as once per iteration with less than 3.1% runtime overhead. Compared to the state-of-the-art methods, LowDiff can shorten training time by up to $89.2\%$ under high checkpointing frequencies (§VI & §VII).
@@ II-A Basics of Distributed Training Systems
To accelerate the training process of large models and datasets [56, 28, 46, 44], distributed training systems are commonly adopted to distribute computations and data across multiple GPUs. Specifically, distributed training has four steps for the $t$-th iteration:
  • Forward pass. The model function is applied to the input data set to obtain the prediction.
  • Backward pass. A loss function is used to calculate the gradient on each worker.
  • Gradient synchronization. The gradients of different workers are synchronized through communication primitives.
  • Model update. The optimizer (e.g., Adam [24]) uses the synchronized gradient $G_{t}$ to update the model state (denoted by $M_{t}$), which contains the model parameters $x_{t}$ and the optimizer parameters $o_{t}$ (i.e., $M_{t}=(x_{t},o_{t})$), satisfying that
$M_{t+1}=M_{t}+Adam(G_{t}).$ (1)
Here, we specially introduce the widely used optimizer, Adam. The Adam optimizer maintains the first-order and second-order moments (i.e., the mean and variance of the gradients), and each has the same size as the model parameters, denoted by $\Psi$. Therefore, Adam requires an additional $2\Psi$ parameter storage beyond the model itself, which will support one of our insights in §III-B.
@@ II-B Checkpointing Techniques
When training large models, failures frequently occur in distributed training systems due to extensive computational resources and long training times. To handle frequent training failures, checkpointing, a widely used technique in common training frameworks (e.g., PyTorch [39] and Tensorflow [1]), has been extensively studied [15, 36, 54, 19, 53, 8, 32]. It regularly copies the model state to local storage or remote storage during training. During recovery from the failure, all GPUs load the latest checkpoint and rerun the remaining process until the point of failure. To evaluate checkpointing performance, recent studies (e.g., [19]) define the wasted time metric as the sum of the recovery time from the latest checkpoint and the steady-state overhead (i.e., the GPU time for checkpointing when no failures have happened).
Frequent checkpointing. Traditional checkpointing is performed at the granularity of epochs, but an epoch may run for hours, thus incurring substantial wasted time for every failure [36]. Thus, state-of-the-art studies [36, 54] that aim for fast recovery increase the checkpointing frequency from epochs to iterations, which we call frequent checkpointing in this paper. The state-of-the-art systems are specified as follows:
Checkfreq [36] decouples checkpointing into snapshot and persist operations, and pipelines them with computation to enable frequent checkpointing. Specifically, the snapshot operation copies model parameters from GPU to CPU memory, while the persist operation asynchronously writes the snapshot to persistent storage. This design allows Checkfreq to create a checkpoint every 14–19 iterations.
Gemini [54] improves frequent checkpointing with the help of the CPU memory of the host machine and proposes a checkpointing traffic scheduling algorithm to mitigate the impact on model training. In this way, Gemini increases the checkpoint frequency over Checkfreq by up to $8\times$, even per-iteration frequency.
PCcheck [52] enables checkpointing as frequently as every 10 iterations by checkpointing to persistent main memory (PMEM). PCcheck utilizes concurrent checkpoints and minimizes time per checkpoint using pipelining techniques and multiple threads.
Just-in-time checkpointing [19] diverges from prior frequent checkpointing techniques. Instead of saving model states at regular intervals, Just-in-time checkpointing initiates a checkpoint only when a failure is detected and assumes redundancy across workers to ensure fault tolerance.
@@
Differential checkpointing. A recent study, Check-N-Run [15], on reducing the checkpointing costs introduces differential checkpointing (DC for short), which records the part of the model that is modified only during the last checkpoint interval so as to reduce the checkpoint size.
Specifically, DC saves its model state as full checkpoint ($C^{F}$) regularly. DC also saves its differential checkpoint of the $t$-th iteration (denoted by $C^{D}_{t}$) at an iteration-level high frequency. Then the model state $M_{n}$ at the $n$-th iteration can be reconstructed by merging the full and all differential checkpoints as
$M_{n}=C^{F}+C^{D}_{1}+C^{D}_{2}+\dots+C^{D}_{n}.$ (2)
However, we note that Check-N-Run implements DC only in recommendation systems, since the recommendation model has a unique property that its embedding table has a high sparsity, such that DC is particularly well suited for recommendation models where only a small part of the model parameters are updated after each iteration.
@@ II-C Gradient Compression Techniques
Gradient compression techniques have been widely used in high-performance distributed training systems [3, 16, 17, 30, 31, 47, 48, 51, 13, 4, 9, 50] to significantly reduce the communication overhead of the gradient synchronization (see §II-A). Two representative compression approaches are Sparsification [3, 51], which selects only a subset of the elements of the gradient, and Quantization [13, 5], which reduces the number of bits of each element of the gradient.
In this paper, we will observe that the costs of deploying DC in distributed training can be largely reduced by leveraging gradient compression (sparsification and quantization), which motivates our main idea (see §III-C for details).
@@ III-A Challenges of DC in General Training
As stated in §II-B, DC may incur high costs when deployed in general distributed training systems without the help of the sparsity feature in recommendation systems, so we conduct measurements to observe that show the challenges of DC in general distributed training systems, especially under high frequency. We conduct our measurement on eight NVIDIA A100 GPUs with a 25Gbps network [54]. We evaluate a typical large-model training task GPT2-L [41], with a common differential checkpointing scheme (see Equation (2)). Detailed setups are shown in §VII-A.
Challenge 1: Frequent DC incurs significant computation cost for highly compressing differentials. In the absence of the sparsity feature’s support, DC has to apply the highly compressed scheme (compression ratio $\rho$ = 0.01 [15]) to the differentials (called compressed differentials), thus degrading the training performance. To measure the impact of compression computation cost on training, Figure 1(a) shows the training time of cases without compression and with different compression frequencies (e.g., 8 iteration compression frequency means the compression is performed every 8 iterations). We see that compared to the case without compression, the case with compression slows down the training process by $13\%\sim 57\%$, while higher frequency incurs slower training.
Challenge 2: Frequent DC causes significant transmission cost due to frequent checkpoint writes. Frequent DC needs to write the differential of model state to storage frequently, which blocks the training process frequently and incurs substantial transmission overhead, since the model update has to wait for the snapshotting of checkpoint [36, 53]. To measure the impact of transmission cost on training in terms of frequency, Figure 1(b) shows the training time of cases without differential transmission and with different differential transmission frequencies. We observe that compared to the case without differential transmission, the case with differential transmission slows down the training process by $12\%\sim 54\%$, while higher frequency incurs slower training.
@@ III-B Findings from Gradient Compression
Finding 1: The compressed gradients can directly construct the compressed differentials for general DNN models. Gradients of a general DNN model are used to adjust the model state via the optimizer (e.g., Adam), so based on Equation (1), we have $Adam(G_{t})=M_{t+1}-M_{t}$. Meanwhile, the differential checkpoint of the $t$-th iteration $C_{t}^{D}$ also equals $M_{t+1}-M_{t}$ by definition (see §II-B). Thus, we have $C_{t}^{D}=Adam(G_{t})$. Therefore, the compressed version of differential checkpoint $C_{t}^{D}$ can be directly constructed via the optimizer by the compressed gradient $\widetilde{G}_{t}$, which can be obtained in advance by the gradient compression operation.

### v2

# Optimizing Frequent Checkpointing via Low-Cost Differential for Distributed Training Systems

@@ Abstract
We proposes LowDiff, an efficient frequent checkpointing framework that reuses compressed gradients, serving as differential checkpoints to reduce cost. Furthermore, LowDiff incorporates a batched gradient write optimization to persist these differentials to storage efficiently. It also dynamically tunes both the checkpoint frequency and the batching size to maximize performance. In non-compression scenario, We further proposes LowDiff+ with a layer-wise gradient reusing and snapshotting approach and a CPU-based asynchronous persistence strategy, enabling frequent checkpointing without gradient compression. Experiments on various workloads show that LowDiff can achieve checkpointing frequency up to per iteration with less than 3.1% runtime overhead.
@@ I Introduction
Deep learning has demonstrated remarkable capabilities in the fields of computer vision [simonyan2014very, he2016deep], natural language processing [devlin2018bert, brown2020language] and so on. Recently, the advent of large models like ChatGPT [chatgpt] and GPT-4 [openai2024gpt4] has captivated both academic and industrial attention. To efficiently train the large models, distributed training systems [li2013parameter, li2014communication, sergeev2018horovod, rasley2020deepspeed] are widely deployed across workers (or nodes) to accelerate the training process, which iterates over a dataset many times (i.e., epochs), each containing multiple iterations. Nevertheless, distributed training often requires many nodes and a long training time, leading to frequent failures [zhang2020empirical]. For instance, a Microsoft study reveals that the mean time between failures (MTBF) in multi-tenant GPU clusters can range from minutes to days [jeon2019analysis]; similarly, OPT-175B, equipped with 992 NVIDIA A100 GPUs, experienced 112 failures during its 90-day training period—averaging nearly two failures per day [OPT].
@@
To handle failures, checkpointing, which regularly writes the model state to storage (e.g., disks) for failure recovery, can confine the potential loss of the training progress to the interval between consecutive checkpoints, only requiring the recovery of the remaining progress. Traditional checkpointing saves the model states (including the model parameters and optimizer parameters) at epoch boundaries, but an epoch often runs for hours, so the epoch-level checkpointing often results in hours of GPU computation wasted on average for every failure recovery [mohan2021checkfreq].
@@
Consequently, state-of-the-art checkpointing studies on fast recovery (e.g., Checkfreq [mohan2021checkfreq] and Gemini [wang2023gemini]) aim to increase the frequency of checkpointing from epochs to iterations (we call frequent checkpointing). Nevertheless, the frequent checkpointing generates a considerable number of checkpoints, whose creation and transmission will incur substantial computational and communication costs, and thus hinder the training process significantly. Recent studies on reducing such costs (e.g., Check-N-Run [eisenman2022check]) introduce differential checkpointing (DC for short), which only tracks and checkpoints the modified part of the model to reduce the checkpoint size, into recommendation systems by leveraging its unique sparse features.
@@
However, when deploying DC with high frequency in general distributed training systems, we observe two challenges arising: (1) without the help of sparse features, frequent DC incurs significant computation cost required for compressing the original differentials into smaller ones (i.e., compressed differentials), thus hindering the training process, and (2) frequent DC causes significant training stalls, as frequent checkpoint writes incur substantial transmissions which can not be overlapped with training, thus degrading the training performance (see §III-A for details).
@@
Fortunately, we also have two findings from the gradient compression techniques [chen2023bbtopk, fei2021efficient, ming2024adtopk, chen2021cswap, song2023optimus] widely used in distributed training systems for communication efficiency: (1) the compressed gradients can be reused to construct the compressed differentials for general deep-learning models. (2) Using compressed gradients eliminates the data dependency between training and checkpointing, enabling them to run in parallel. Moreover, the size of compressed gradients is significantly smaller than that of compressed differentials (see §III-B for details).
@@
Therefore, to address the above challenges, the corresponding findings inspire us to leverage the compressed gradients in distributed training systems for low-cost differential checkpointing, so as to improve the overall training performance. Our motivation is that the compressed gradients can act as the compressed differentials with no DC computation cost and with a lower DC transmission cost. Based on the findings, we propose LowDiff, an efficient frequent checkpointing framework, whose main idea is to reuse the compressed gradients to serve as differential checkpoints, which not only eliminates the computation cost for differential compression (via reusing) but also reduces the transmission cost for DC writes (due to smaller-size gradients than differentials). Further, to enhance LowDiff, we accelerate the checkpoint writes via gradient batching and optimize the checkpointing performance by tuning the checkpoint frequency and batching size.
@@
  • We perform an analysis of LowDiff, showing that gradient reuse eliminates the data dependency and enables parallel execution between checkpointing and training. Through measurement, we further demonstrate that checkpointing can be overlapped with the training iteration, allowing checkpointing to execute concurrently with training (§IV).
  • We design LowDiff, which realizes the reuse of compressed gradients via a queue structure and facilitates zero-copy transmission between the training and checkpointing processes. Atop LowDiff, we propose a batched gradient writing optimization by merging compressed gradients in CPU memory before persisting to storage. We also design a checkpointing configuration optimization to minimize the wasted GPU time (§V).
  • We further propose LowDiff+ for efficient gradient reuse without compression. By exploiting the layer-by-layer generation of gradients during training, we design a layer-wise gradient reusing and snapshotting approach. We also reconstruct a model state in CPU memory, performing high-frequency in-memory checkpointing and asynchronous persistence (§VI).
  • We implement LowDiff and LowDiff+ atop DeepSpeed, open-sourced at https://github.com/YuchongHu/LowDiff. Experiments on various models show that LowDiff and LowDiff+ can achieve checkpointing as frequently as once per iteration with less than minimal runtime overhead. Compared to the state-of-the-art methods, LowDiff and LowDiff+ can shorten training time by up to $89.2\%$ and $81.7\%$ under high checkpointing frequencies (§VII & §VIII).
@@ II-A Basics of Distributed Training Systems
To accelerate the training process of large models and datasets [zhao2023pytorch, li2020pytorch, sergeev2018horovod, rasley2020deepspeed], distributed training systems are commonly adopted to distribute computations and data across multiple GPUs. Specifically, distributed training has four steps for the $t$-th iteration:
  • Forward pass. The model parameters $x_{t}$ is applied to the input data set $X$ to obtain the prediction $Y$ as:
    $Y=Forward(x_{t},X).$ (1)
  • Backward pass. A loss function is used to calculate the gradient $G_{i,t}$ on $i$-th worker, satisfying that
    $G_{i,t}=Backward(x_{t},Y).$ (2)
    Then the gradients of workers are synchronized through communication primitives, satisfying that
    $G_{t}=Sync(G_{i,t}).$ (3)
  • Model update. The optimizer (e.g., Adam [kingma2014adam]) uses the synchronized gradient $G_{t}$ to update the model state (denoted by $M_{t}$), which contains the model parameters $x_{t}$ and the optimizer parameters $o_{t}$ (i.e., $M_{t}=(x_{t},o_{t})$), satisfying that
    $M_{t+1}\leftarrow M_{t}+Adam(G_{t}).$ (4)
Here, we specially introduce the widely used optimizer, Adam. The Adam optimizer maintains the first-order and second-order moments (i.e., the mean and variance of the gradients), and each has the same size as the model parameters, denoted by $\Psi$. Therefore, Adam requires an additional $2\Psi$ parameter storage beyond the model itself, which will support one of our findings in §III-B.
@@ II-B Checkpointing Techniques
When training large models, failures frequently occur in distributed training systems due to extensive computational resources and long training times. To handle frequent training failures, checkpointing, a widely used technique in common training frameworks (e.g., PyTorch [paszke2019pytorch] and Tensorflow [abadi2016tensorflow]), has been extensively studied [eisenman2022check, mohan2021checkfreq, wang2023gemini, gupta2024just, wang2024fastpersist, chen2023cost, maurya2024datastates]. It regularly copies the model state to local storage or remote storage during training. During recovery from the failure, all GPUs load the latest checkpoint and rerun the remaining process until the point of failure. To evaluate checkpointing performance, recent studies (e.g., [gupta2024just]) define the wasted time metric as the sum of the recovery time from the latest checkpoint and the steady-state overhead (i.e., the GPU time for checkpointing when no failures have happened).
Frequent checkpointing. Traditional checkpointing is performed at the granularity of epochs, but an epoch may run for hours, thus incurring substantial wasted time for every failure [mohan2021checkfreq]. Thus, state-of-the-art studies [mohan2021checkfreq, wang2023gemini] that aim for fast recovery increase the checkpointing frequency from epochs to iterations, which we call frequent checkpointing in this paper. The state-of-the-art systems are specified as follows:
Checkfreq [mohan2021checkfreq] decouples checkpointing into snapshot and persist operations, and pipelines them with computation to enable frequent checkpointing. Specifically, the snapshot operation copies model parameters from GPU to CPU memory, while the persist operation asynchronously writes the snapshot to persistent storage. This design allows Checkfreq to create a checkpoint every 14–19 iterations.
Gemini [wang2023gemini] improves frequent checkpointing with the help of the CPU memory of the host machine and proposes a checkpointing traffic scheduling algorithm to mitigate the impact on model training. In this way, Gemini increases the checkpoint frequency over Checkfreq by up to $8\times$, even per-iteration frequency.
PCcheck [strati2025pccheck] enables checkpointing as frequently as every 10 iterations by checkpointing to persistent main memory (PMEM). PCcheck utilizes concurrent checkpoints and minimizes time per checkpoint using pipelining techniques and multiple threads.
Just-in-time checkpointing [gupta2024just] diverges from prior frequent checkpointing techniques. Instead of saving model states at regular intervals, Just-in-time checkpointing initiates a checkpoint only when a failure is detected and assumes redundancy across workers to ensure fault tolerance.
@@
Differential checkpointing. A recent study, Check-N-Run [eisenman2022check], on reducing the checkpointing costs introduces differential checkpointing (DC for short), which records the part of the model that is modified only during the last checkpoint interval so as to reduce the checkpoint size.
Specifically, DC saves its model state as full checkpoint ($C^{F}$) regularly. DC also saves its differential checkpoint of the $t$-th iteration (denoted by $C^{D}_{t}$) at an iteration-level high frequency as
$C^{D}_{t}=M_{t+1}-M_{t}.$ (5)
Then the model state $M_{t}$ at the $t$-th iteration can be reconstructed by merging the full and all differential checkpoints as
$M_{t}=C^{F}+C^{D}_{1}+C^{D}_{2}+\dots+C^{D}_{t-1}.$ (6)
However, we note that Check-N-Run implements DC only in recommendation systems, since the recommendation model has a unique property that its embedding table has a high sparsity such that DC is particularly well suited for recommendation models where only a small part of the model parameters are updated after each iteration.
@@ II-C Gradient Compression Techniques
Gradient compression techniques have been widely used in high-performance distributed training systems [aji2017sparse, faghri2020adaptive, fang2019redsync, lin2017deep, m2021efficient, shi2019layer, shi2019distributed, stich2018sparsified, dettmers20158, alistarh2017qsgd, chen2021cswap, song2023optimus] to significantly reduce the communication overhead of the gradient synchronization in backward pass (see §II-A and Equation (3)). Two representative compression approaches are Sparsification [aji2017sparse, stich2018sparsified], which selects only a subset of the elements of the gradient, and Quantization [dettmers20158, NIPS2017QSGD], which reduces the number of bits of each element of the gradient.
In this paper, we will observe that the costs of deploying DC in distributed training can be largely reduced by leveraging gradient compression, which motivates our main idea (see §III-C for details).
@@ III-A Challenges of DC in General Training
As stated in §II-B, DC may incur high costs when deployed in general distributed training systems without the help of the sparsity feature in recommendation systems, so we conduct measurements to observe that show the challenges of DC in general distributed training systems, especially under high frequency. We conduct our measurement on eight NVIDIA A100 GPUs with a 25Gbps network [wang2023gemini]. We evaluate a typical large-model training task GPT2-L [radford2019language], with a common differential checkpointing scheme (see Equation (6)). Detailed setups are shown in §VIII-A.
Challenge 1: Frequent DC incurs significant computation cost for highly compressing differentials. In the absence of the sparsity feature’s support, DC has to apply the highly compressed scheme (compression ratio $\rho$ = 0.01 [eisenman2022check]) to the differentials (called compressed differentials), thus degrading the training performance. To measure the impact of compression computation cost on training, Figure 1(a) shows the training time of cases without compression and with different compression frequencies (e.g., 8 iteration compression frequency means the compression is performed every 8 iterations). We see that compared to the case without compression, the case with compression slows down the training process greatly by $13\%\sim 57\%$, while higher frequency incurs much slower training.
Challenge 2: Frequent DC causes significant transmission cost due to frequent checkpoint writes. Frequent DC needs to write the differential of model state to storage frequently, which blocks the training process frequently and incurs substantial transmission overhead, since the model update has to wait for the snapshotting of checkpoint [mohan2021checkfreq, wang2024fastpersist]. To measure the impact of transmission cost on training in terms of frequency, Figure 1(b) shows the training time of cases without differential transmission and with different differential transmission frequencies. We observe that compared to the case without differential transmission, the case with differential transmission slows down the training process by $12\%\sim 54\%$, while higher frequency incurs slower training.
@@ III-B Findings from Gradient Compression
Finding 1: The compressed gradients can directly construct the compressed differentials for general DNN models. Gradients of a general DNN model are used to adjust the model state via the optimizer (e.g., Adam), so based on Equation (4), we have $Adam(G_{t})=M_{t+1}-M_{t}$. Meanwhile, the differential checkpoint of the $t$-th iteration $C_{t}^{D}=M_{t+1}-M_{t}$ based on Equation (5). Thus, we have:
$C_{t}^{D}=Adam(G_{t}).$ (7)
Therefore, the compressed version of the differential checkpoint $C_{t}^{D}$ can be directly constructed via the optimizer by the compressed gradient $\widetilde{G}_{t}$, which can be obtained in advance by the gradient compression operation.